OpenStudy (anonymous):

(11^-4)^-6=11 "Enter the answer." This is how the question was written, if anyone gives correct information and explains it for me, they will receive a medal

imqwerty (imqwerty):

it must be like this ig-\[(11^\left( -4 \right))^\left( -6 \right)=11^x\] ?

imqwerty (imqwerty):

if that is the case then we can apply this law to get the answer- \[(x^a)^b =x^\left( a \times b \right)\]

imqwerty (imqwerty):

so we have- \[(11^\left( -4) \right))^\left( -6 \right) =11^\left( -4 \times -6 \right)\] \[=11^\left( 24 \right)\] so \[11^\left( 24 \right)=11^x\] x=24

OpenStudy (anonymous):

ty @imqwerty